OpenStudy (luigi0210):

Find the area:

OpenStudy (luigi0210):

\[\large \int\limits_{0}^{\pi} sinx~dx\]

OpenStudy (psymon):

Bored?

OpenStudy (luigi0210):

And reviewing ;P

terenzreignz (terenzreignz):

Antiderivative of sin(x) is...?

OpenStudy (psymon):

Lol, okay then. Integral sinx? We remember that?

OpenStudy (luigi0210):

Uh, -cosx?

OpenStudy (inkyvoyd):

damnit this problem is easy D:

terenzreignz (terenzreignz):

Well, if you're that bored, you might as well use \[\Large e^{2\pi^3}-\cos(x)\]

OpenStudy (psymon):

Lol.

OpenStudy (psymon):

\[\int\limits_{}^{}e ^{x^{2}}dx\]

terenzreignz (terenzreignz):

Well.... there's nothing we can do about that ^ Unless you're suggesting series...

OpenStudy (luigi0210):

Psymon, what's a double integral?

terenzreignz (terenzreignz):

This: \[\Large \iint\limits_R f(x,y)dA\]? :D

OpenStudy (psymon):

Lol, actually, there is something you can do about the one I posted. Thats why I specifically posted that one.

terenzreignz (terenzreignz):

None that I know of. :/

OpenStudy (psymon):

Its actually a special function. Its called an error function, lol. And then theres also a complementary error function.

OpenStudy (psymon):

Theres all these really weird special functions Im being shown in class :/

OpenStudy (psymon):

\[F(\pi) - F(0)\] \[-\cos(\pi) - [-\cos(0)]\] @Luigi0210

terenzreignz (terenzreignz):

@Psymon You forgot something :3 \[\Large \color{green}{e^{2\pi^3}}-\cos(\pi) - [\color{green}{e^{2\pi^3}}-\cos(0]\]

OpenStudy (psymon):

Not in reference to yours D: @terenzreignz

OpenStudy (luigi0210):

Hey Psymon is this right? \[\int\limits_{0}^{1}(\int\limits_{0}^{2}xy^2~dx)dy\] \[\int\limits_{0}^{1}(2y^2)dy=\frac{2}{3}\]

OpenStudy (psymon):

Yep. I think so at first glance.

OpenStudy (psymon):

Didnt do any paper, though.

OpenStudy (luigi0210):

I barely know how to do single integrals o_O

OpenStudy (psymon):

Double and triple integrals are easy. Same with partial derivatives. Its only when you have to start doing conversions an dsubstitutions with the integrals that things get interesting.

OpenStudy (luigi0210):

Well for you, I have very little experience with integrals

OpenStudy (psymon):

You did it right, though O.o

terenzreignz (terenzreignz):

\[\int\limits_{0}^{1}\int\limits_{0}^{2}(xy^2)~dxdy=\int\limits_0^1x~dx\int\limits_0^2y^2dy\]

OpenStudy (psymon):

Is it really back integral first? O.o

terenzreignz (terenzreignz):

Doesn't matter since the x and y expressions are separable from each other.

OpenStudy (psymon):

Okay, yeah, i havent ever worked with doubles and triples, I just know the concept xD So you show me something and its likely to benew, haha.

terenzreignz (terenzreignz):

Well, there's something wrong \[\Large \frac12\cdot \frac{8}3\]

terenzreignz (terenzreignz):

\[\Large = \frac43 \ne \frac23\]

OpenStudy (psymon):

Eh O.o

OpenStudy (psymon):

Well, I know what luigi did to get his answer, so what did you do.

terenzreignz (terenzreignz):

And what do you get when you integrate y^2? -.-

OpenStudy (luigi0210):

Sorry I haven't slept xD

terenzreignz (terenzreignz):

Unless... I got it reversed

OpenStudy (psymon):

Okay, so seriously, why is it multiplication instead of what im thinking is going on? \[\int\limits_{0}^{1}\int\limits_{0}^{2}xy^{2}dxdy\] \[\frac{ x^{2} y^2}{ 2 }\] F(2) - F(0) = 2y^2 \[\int\limits_{0}^{1}2y^{2}dy \] \[\frac{ 2y^{3} }{ 3 } \] F(1) - F(0) = 2/3 What am I not getting?

terenzreignz (terenzreignz):

My mistake :D

terenzreignz (terenzreignz):

Sorry, must have not integrated with the correct limits... must have switched them

terenzreignz (terenzreignz):

hehe... whoops

OpenStudy (psymon):

Lol, no biggie xD

OpenStudy (luigi0210):

._.

OpenStudy (psymon):

Luigi be correcti :3

OpenStudy (luigi0210):

I'm going back to algebra >.>

OpenStudy (psymon):

Hey, you got it, no need to xD Why the double integral question, though?

terenzreignz (terenzreignz):

I'm logging out... I've caused enough trouble for today :3 Catch you guys later ^_^ ------------------------------------ TJ out

OpenStudy (luigi0210):

I was curious? Idk -_- See yall TJ Now.. how do you graph this? >.<

OpenStudy (psymon):

xy^2?

OpenStudy (luigi0210):

Yea, whatever that means :/

OpenStudy (psymon):

Im not sure really. It doesnt fit any of the forms I would know how to graph, haha x_x

OpenStudy (psymon):

I mean, I can kinda guess.

OpenStudy (luigi0210):

\[z=xy^2\] \[0 \le x \le2 \] \[0\le x \le 1\] My calculuator gave me this >.>

OpenStudy (luigi0210):

*\[0 \le y \le 1\]

OpenStudy (psymon):

I couldnt tell ya what my graphing calculator got. TOo hard to perfectly tell, lol.

OpenStudy (luigi0210):

I need to go back to watching videos >.< Thanks Psymon :)

OpenStudy (psymon):

Alright, laterz xD